John is 4 years older than his sister and 6 years ago he was twice his sister's age then. How old is each now?
Answer by ankor@dixie-net.com(22740)

Let j = John's age now
let s = sister's age now
:
Write an equation for each statement:
:
"John is 4 years older than his sister"
j = s + 4
:
"6 years ago he was twice his sister's age then."
(j-6) = 2(s-6)
j - 6 = 2s - 12
j = 2s - 12 + 6
j = 2s - 6
:
How old is each now?
:
From the 1st equation, substitute (s+4) for j in the last equation
s + 4 = 2s - 6
4 + 6 = 2s - s
10 = s (sister's age now)
then
10 + 4 = 14 yr John's age now
:
:
Check solution in the statement:
"6 years ago he was twice his sister's age then."
14 - 6 = 2(10 - 6)